OpenStudy (anonymous):

How do I solve the ration equation 1 / 2 + 1 / 3 = 1 / x?

OpenStudy (anonymous):

@drew123 do you have any possible answers?

OpenStudy (anonymous):

Nope, I had to make the equation myself and now solve and show work

OpenStudy (anonymous):

@drew123 oh ok one sec

OpenStudy (anonymous):

1 / 2 + 1 / 3 = 1 / x? make them have a common denominator 3/6 + 2/6 = 5/6 = 1/(6/5)

OpenStudy (anonymous):

Is that the answer, or just a step?

OpenStudy (anonymous):

well, I think that that would be how to solve for 1/x

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're welcome